Divide Rs 1,260 among A,B and C so that the ratio between the shares of A and B is 2:3 and the ratio between the shares of B and C is 4:5.pls.. give me the solution.
Answer by fcabanski(1391)

The problem gives the following:


A+B+C=1260
A/B = 2/3 cross multiply 3A = 2B and B = 3/2 * A
B/C = 4/5 cross multiply 5B = 4C and C = 5/4 * B


Solve the first equation for A by substituting in B and C in terms of A.


B = 3/2 * A and C = 5/4 * B = 5/4 * 3/2 * A = 15/8 * A


A + (3/2)A + (15/8)A = 1260
(8A + 12A + 15A)/8 = 1260
35A = 10080
A=288


B = 3/2 * A = 3/2 * 288 = 432
C = 5/4 * B = 5/4 * 432 = 540

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com